Assume that the average firm in your company's industry is expected to grow at a constant rate of 4% and that its dividend yield is 8%. Your company is about as risky as the average firm in the industry and just paid a dividend (D0) of $2.75. You expect that the growth rate of dividends will be 50% during the first year (g0,1 = 50%) and 25% during the second year (g1,2 = 25%). After Year 2, dividend growth will be constant at 4%. What is the estimated value per share of your firm’s stock? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required rate=Dividend yield+Growth rate

=4+8=12%

D1=(2.75*1.5)=4.125

D2=(4.125*1.25)=5.15625

Value after year 2=(D2*Growth rate)/(Required rate-Growth rate)

=(5.15625*1.04)/(0.12-0.04)

=67.03125

Hence estimated value=Future dividend and value*Present value of discounting factor(rate%,time period)

=4.125/1.12+5.15625/1.12^2+67.03125/1.12^2

=$61.23(Approx).